if 1/6sinx , cosx , tanx, are in g.p , the value of x is equal to ?

Grade:11

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear honey,

If they are in GP,then

cos2x=sinx/6sinxcosx

cos3x=1/6

x=cos-1(1/cuberoot of 6)

We are all IITians and here to help you in your IIT JEE preparation.

All the best.

 If you like this answer please approve it....

win exciting gifts by answering the questions on Discussion Forum

Sagar Singh

B.Tech IIT Delhi